# Jingyao Liu♠♣  Chen Huang♠♢†  Zhizhao GuanWenqiang Lei♠♣Yang Deng

### v1

@@ 4 Experiment
Table 3: Benchmark analysis across various LLM backbones regarding the effectiveness and efficiency. Vanilla LLM is highlighted in light grey as a reference. For effectiveness, the highest effective metric value within each group is shown in bold, and the second highest is underlined.
Req. Acc. Test Acc Balance Score Cost (USD) Footprint (gCO2eq) Duration (s)
[…]
GPT-Engineer 50.37 66.68 56.90 0.0198 0.132 21
Self-Collaboration 46.14 60.58 51.92 0.0155 0.109 37
MapCoder 47.92 63.89 54.31 0.1091 0.750 93
ChatDev 42.75 58.08 48.88 0.1947 1.910 114
MetaGPT 0.00 0.18 0.07 0.0951 0.794 66
[…]
GPT-Engineer 43.10 58.45 49.24 0.0012 0.005 18
Self-Collaboration 38.76 53.98 44.85 0.0009 0.004 25
MapCoder 40.89 57.41 47.50 0.0072 0.033 88
ChatDev 33.95 49.85 40.31 0.0118 0.071 157
MetaGPT 0.00 0.27 0.11 0.0067 0.040 63
[…]
GPT-Engineer 49.79 64.13 55.52 0.0030 0.067 66
Self-Collaboration 42.68 59.22 49.30 0.0022 0.053 75
MapCoder 48.37 64.25 54.72 0.0186 0.424 366
ChatDev 43.07 58.95 49.43 0.0249 0.790 300
MetaGPT 1.63 2.84 2.11 0.0207 0.544 312
[…]
GPT-Engineer 42.08 58.78 48.76 0.0037 0.080 45
Self-Collaboration 42.57 55.57 47.77 0.0030 0.066 69
MapCoder 40.44 57.53 47.28 0.0293 0.624 294
ChatDev 43.52 57.67 49.18 0.0387 1.006 341
MetaGPT 0.00 0.17 0.07 0.0290 0.786 151
[…]
GPT-Engineer 24.03 39.94 30.39 0.0003 0.007 31
Self-Collaboration 20.58 33.55 25.76 0.0003 0.006 52
MapCoder 12.83 27.65 18.76 0.0024 0.049 236
ChatDev 11.30 20.43 14.95 0.0075 0.187 344
MetaGPT 0.00 1.83 0.43 0.0207 0.095 301
@@ 4.1 Experiment Setup
Baseline & LLM Backbones. We select a diverse set of representative E2ESD frameworks across various LLM backbones. These include: (1) Vanilla LLM that generates the full codebase via simply prompting LLM using user requirements without external scaffolding. (2) Single-Agent Framework, GPT-Engineer (Engineer, 2024), which builds the complete codebase in a single reasoning pass with minimal external coordination while maintaining a modular development workflow. (3) Multi-Agent Frameworks, a dominant paradigm, where the development process is decomposed into specialized agent roles based on classic software engineering principles (Sommerville, 2011). We benchmark with a wide range of state-of-the-art multi-agent framework, including Self-Collaboration (Dong et al., 2024), MetaGPT (Hong et al., 2024), MapCoder (Islam et al., 2024), and ChatDev (Qian et al., 2024).
Finally, we adopt a range of backbones from two vendors (Hurst et al., 2024; Hui et al., 2024), covering various scales and architectures: GPT-4o, GPT-4o-mini, Qwen-7B (Qwen2.5-7B-Instruct), Qwen-70B (Qwen2.5-72B-Instruct), and the mixture-of-experts (MoE) Qwen-Max (Qwen2.5-Max). This enables controlled analysis of framework versus model backbone effects. Implementation details are presented in Appendix B.1.
Evaluation Metrics. We propose a systematic evaluation protocol for E2ESD, involving both code effectiveness (how well the generated code meets user requirements) and generation efficiency (how is the cost and time of the code generation process). As for code effectiveness, the evaluation metrics include: (1) Req. Acc, the proportion of requirements successfully validated by all their corresponding test cases across a project, which measures the requirement-level accuracy; (2) Test Acc, the proportion of all passed test cases across a project, which measures the test-level accuracy; (3) Balanced Score, a weighted combination of Req. Acc and Test Acc to mitigate bias from varying test case granularity per requirement. Additionally, generation efficiency is assessed using the following three metrics: (1) Cost (USD), the average API pricing for LLM token usage; (2) Carbon Footprint (CO2), the average carbon footprint for generating one project; (3) Duration, the average wall-clock time for project generation. Metric details are presented in Appendix B.2.
@@ 4.2 Main Results
How effective are existing methods on E2ESD? Despite meeting broad project requirements, performance remains below acceptable standards due to lack of proficiency in handling specifics. Off-the-shelf models excel at function-level tasks (e.g., HumanEval (Chen et al., 2021) in Appendix C.3) but underperform on project-level E2ESD, achieving only 30%–50% Req. Acc. on the proposed E2EDev dataset (Table 3), with even GPT-4o below 60%. Furthermore, Test Acc. scores, which allow partial fulfillment, suggest failures stem from their lack of proficiency in handling detailed functional specifics during implementation. This phenomenon, visualized in Figure 4, persists across different LLM backbones and E2ESD frameworks. The observed discrepancy of over 25% in both metrics suggests that while models can implement the required functionality, they often fail to address complex edge cases. Importantly, this issue is also reflected in $\sim$10% performance gap observed on HumanEval and HumanEval-ET (cf. Appendix C.3).
@@
Figure 4: Soft Req. Acc. and Req. Acc. under three representative LLMs. Here, V-LLM refers to the Vanilla LLM, GPT-E refers to GPT-Engineer, and Self-C refers to Self-Collaboration. Additional results on other backbones are reported in Appendix C.1, Table 8.
How efficient are existing methods on E2ESD? Streamlined workflows offer high efficiency, but multi-agent frameworks often exhibit excessive interaction rounds and token costs with minimal effectiveness gains. As evidenced by the efficiency-oriented metrics in Table 3, methods such as Vanilla LLM, GPT-Engineer, and Self-Collaboration maintain high efficiency ($\leq$ three times API Calls), benefiting from their streamlined, deterministic workflows. However, regarding the multi-agent frameworks, even static frameworks like MapCoder and MetaGPT require >10 interaction turns per task, amplifying latency and computational expense. Dynamic frameworks (e.g., ChatDev with GPT-4o) reach 15.72 turns on average (minimum of 9), with inefficiency exacerbated by repetitive and uninformative dialogue cycles, as analyzed in detail in Appendix C.2.
Figure 5: Comparisons across agentic frameworks and the Vanilla LLM (on Req. Acc.). Bars within the gray zone indicate lower performance.
Figure 6: Statistics of interaction overhead across representative frameworks. P. Tok., C. Tok., and Turns refer to average prompt tokens, completion tokens, and dialogue rounds, respectively.
Model Method P.Tok C.Tok Turns
@@
Does the integration of agentic frameworks consistently enhance the effectiveness of the Vanilla LLM? Not necessarily. While agentic frameworks offer the potential for performance gains, their inherent complexity imposes significant demands on the foundational capabilities of the underlying LLM. Figure 6 shows framework performance varies widely across LLMs, with Vanilla LLMs occasionally surpassing frameworks (and vice versa), revealing architectural overhead. Notably, multi-agent frameworks often underperform single-agent approaches due to the necessary coordination demands among agents, such as task decomposition (Xia et al., 2024; Han et al., 2024), role/instruction following (Pan et al., 2025; Hammond et al., 2025), dialogue state/context management (Pan et al., 2025; Li et al., 2023; Hammond et al., 2025), and the increasing interaction costs as shown in Table 6. Therefore, error accumulation from these complexities can degrade performance, particularly in E2ESD tasks. We posit that current frameworks tend to rely heavily on LLMs with robust and versatile foundational capabilities. This suggests that future framework designs should prioritize minimizing this reliance to enhance their practical utility and robustness across diverse LLM backbones.
Why is the performance of MetaGPT notably low on E2ESD? It suffers from communication breakdowns within its specialized multi-agent architecture. While confirming MetaGPT’s reported function-level performance on HumanEval using the author’s code (cf. Appendix C.3), MetaGPT fails to handle nearly all test cases and requirements at the project level, even when using powerful LLMs like GPT-4o or Qwen-Max. This issue seems to be communication breakdowns within this multi-agent framework, ultimately undermining its efficacy (further discussions in Section 4.3).
4.3 Understanding the Limitations of Existing Frameworks on E2ESD
Setup. To gain deeper insights, we conduct a rigorous human evaluation. To balance analytical depth with the practical constraints of human effort, we randomly select 10 data entries from our dataset, resulting in 300 generated projects. Four domain experts then interact with these generated projects, guided by the corresponding user requirements, and evaluate across four key aspects detailed in Table 4 (see Appendix B.3 for evaluation protocol). Finally, Figure 7 visualizes the failure mode distributions of existing methods that employ representative LLMs as backbones — known to be effective in agentic frameworks — grouped by framework type, with detailed results and case studies for each framework provided in the Appendix C.1 and Appendix C.4, respectively.
Table 4: Classification of errors in LLM-generated projects
Error Type Description
Code Inconsistency Internal inconsistencies in generated code, such as unimplemented functions, conflicting or duplicated implementations, empty functions, or signature mismatches.
Requirement Missing Required functionality is not implemented.
Requirement Misaligned Logic of implemented functionality deviates from requirements.
Detail Mismatch Functionality logic aligns with requirements, but contains inaccurate details.
Figure 7: Requirement-level error distribution across different frameworks. Pie segments show error proportions; the center shows total errors across three representative LLM backbones. Trend lines below show the absolute values of each specific error across different frameworks.
Analysis of Code Inconsistency While Vanilla LLMs maintain high code consistency, integrating these LLMs with the E2ESD framework may compromise this consistency, particularly within multi-agent frameworks. This is likely due to the potential exposure to excessive or irrelevant information within the context and prompt. For example, MapCoder employs analogical prompting (Yasunaga et al., 2024), in which the LLM first retrieves project exemplars from its own memory (i.e., generate exemplars by itself), and then leverages them for few-shot enhancement before tackling the given project. Although this approach is intended to provide richer contextual support, our analysis reveals that these retrieved projects often have only a loose relationship to the current task, sharing a general domain (e.g., web applications) but lacking significant functional similarity. This extraneous information introduces noise rather than clarity. In contrast, frameworks that restrict the coder agent’s input to task-specific requirements and component analyses minimize interference, resulting in more consistent code generation.
Analysis of Requirement Missing Flawed multi-agent architecture design may contribute to overlooked requirements. As shown in Figure 7, Vanilla LLM and ChatDev, the only frameworks failing to incorporate requirement analysis during code generation [^4] GPT-Engineer implicitly performs requirement analysis before code generation via “think step by step”., exhibit severe Missing Requirement issue. While ChatDev assigns an agent for requirement analysis, its architecture limits inter-agent communication to pairwise interactions. Consequently, its coding agent interacts solely with the Executive Officer, receiving only high-level structural guidance (e.g., HTML structure, JavaScript functionality, CSS styling) and lacking detailed requirement specifications. This limited visibility significantly elevates the risk of overlooking critical requirements. Thus, ensuring coding agents have access to comprehensive requirement analysis is essential for complete software solutions.
@@ 4.3 Understanding the Limitations of Existing Frameworks on E2ESD
Analysis of Requirement Misaligned Core component analysis helps align code with user requirements. Figure 7 shows multi-agent frameworks significantly reduce Requirement Misaligned issues compared to others. This improvement likely stems from explicitly integrating core component analysis—specifically, the identification of key structural and functional elements such as HTML layout, JavaScript logic (script.js), and CSS styles (styles.css). This analysis informs the coding agent, allowing a focus on user-driven logic and structure dictated by user requirements. Consequently, the generated code better reflects intended functionality and aligns with user requirements.
Analysis of Detail Mismatch LLMs inherently struggle with fine-grained detail control, a limitation amplified by the additional context introduced by multi-agent frameworks. Trend lines in Figure 7 reveal that detail mismatches are prevalent across all frameworks, with multi-agent frameworks moderately increasing their frequency—Vanilla LLM produces the fewest (47), while Self-Collaboration generates the most (68). To gain more sights, we categorize detail mismatches into three types: (1) unhandled logical edge cases, where interdependent features lead to state synchronization failures; (2) missing error handling or validation logic; and (3) UI component display inconsistencies. In particular, logical edge cases remain particularly challenging; for instance, in a dictionary-based favorite-word feature, models consistently fail to update the “favorited” status after removing a word. Missing error handling typically results from absent conditional branches for exceptional cases, and multi-agent frameworks often overlooks these despite explicit instructions in the original prompt—likely due to omission during requirement analysis or component planning. To mitigate this, frameworks like MapCoder and ChatDev prepend each agent’s input with both the original user requirements and the previous agent’s message, preserving contextual fidelity and reducing omissions. Finally, UI display mismatches, such as incorrect decimal precision or text rendering, may largely attribute to LLMs’ intrinsic tendencies, and persist even under strict prompting.
@@
Detailed Analysis on MetaGPT Refer to caption Figure 8: Error distribution of MetaGPT Communication breakdowns within its multi-agent architecture significantly impair code consistency and requirement fidelity. As seen in Figure 8, 44% of all failures stem from code consistency issues, like missing files and syntax errors. Notably, over 43% of these are caused by the programmer agent ignoring the architect’s file structure, despite explicit constraints. Additionally, 54% of syntax errors involve malformed imports or unmatched tokens, often resulting from the engineer agent being tasked with both code generation and tool invocation—prompt instructions for the latter outweigh the former by over 10×, introducing noise that disrupts code generation. Furthermore, 25% of failures are due to missing requirements, with another 31% stemming from inconsistent or incomplete implementation. This may stem from the product manager frequently rewriting or compressing the original user requirement, despite being instructed to restate it verbatim. We observe similar trends in HumanEval: MetaGPT with GPT-4o only achieves $\sim$50% pass rate, with $\sim$30% of completions failing to match required function names. This suggests that an over-engineered agent workflow degrades both adherence to specification and coding reliability.
5 Conclusion & Discussion

### v4

@@ 4 Experiment
Backbone Method Effectiveness ($\uparrow$, %) Efficiency ($\downarrow$)
Req. Acc. Test Acc Bal. Score Cost (\$) CO2 (g) Time (s)
Claude-Haiku 4.5 Vanilla LLM 48.69 $\pm$1.30 63.08 $\pm$0.50 54.45 $\pm$0.98 0.015 0.028 53
GPT-Engineer 53.75 $\pm$0.43 69.41 $\pm$1.07 60.01 $\pm$0.69 0.016 0.038 59
Self-Collab. 49.01 $\pm$1.74 61.50 $\pm$1.26 54.01 $\pm$1.55 0.014 0.033 75
MapCoder 49.61 $\pm$0.50 65.65 $\pm$0.44 56.03 $\pm$0.48 0.099 0.081 129
ChatDev 44.73 $\pm$1.50 58.09 $\pm$1.03 50.07 $\pm$1.31 0.174 0.188 335
MetaGPT 5.39 $\pm$0.76 10.66 $\pm$1.77 7.50 $\pm$1.16 0.090 0.091 366
GPT-4o Vanilla LLM 45.95 $\pm$1.25 60.88 $\pm$1.97 51.92 $\pm$1.54 0.0160 0.083 28
GPT-Engineer 50.83 $\pm$1.33 66.59 $\pm$2.43 57.13 $\pm$1.77 0.0198 0.132 21
Self-Collab. 46.83 $\pm$1.12 61.15 $\pm$1.47 52.56 $\pm$1.26 0.0155 0.109 37
MapCoder 47.70 $\pm$2.57 63.97 $\pm$2.25 54.21 $\pm$2.44 0.1091 0.750 93
ChatDev 42.71 $\pm$2.44 58.93 $\pm$2.66 49.20 $\pm$2.53 0.1947 1.910 114
MetaGPT 0.00 $\pm$0.00 0.17 $\pm$0.05 0.07 $\pm$0.02 0.0951 0.794 66
GPT-4o-mini Vanilla LLM 44.82 $\pm$2.01 60.65 $\pm$2.49 51.15 $\pm$2.20 0.0010 0.003 16
GPT-Engineer 42.13 $\pm$1.81 57.51 $\pm$1.98 48.28 $\pm$1.88 0.0012 0.005 18
Self-Collab. 37.90 $\pm$1.92 52.99 $\pm$2.51 43.94 $\pm$2.16 0.0009 0.004 25
MapCoder 41.30 $\pm$1.49 57.87 $\pm$0.85 47.93 $\pm$1.23 0.0072 0.033 88
ChatDev 33.16 $\pm$2.71 48.40 $\pm$2.75 39.26 $\pm$2.73 0.0118 0.071 157
MetaGPT 0.00 $\pm$0.00 0.23 $\pm$0.05 0.09 $\pm$0.02 0.0067 0.040 63
Qwen-Max Vanilla LLM 43.33 $\pm$0.66 58.57 $\pm$0.73 49.43 $\pm$0.69 0.0025 0.043 53
GPT-Engineer 49.61 $\pm$0.60 63.70 $\pm$1.41 55.25 $\pm$0.92 0.0030 0.067 66
Self-Collab. 42.30 $\pm$1.88 60.44 $\pm$1.33 49.56 $\pm$1.66 0.0022 0.053 75
MapCoder 48.83 $\pm$1.68 64.00 $\pm$1.82 54.90 $\pm$1.74 0.0186 0.424 366
ChatDev 43.93 $\pm$1.94 59.44 $\pm$0.99 50.13 $\pm$1.56 0.0249 0.790 300
MetaGPT 1.65 $\pm$0.25 2.80 $\pm$0.58 2.11 $\pm$0.38 0.0207 0.544 312
Qwen-70B Vanilla LLM 35.75 $\pm$0.00 53.14 $\pm$0.00 42.71 $\pm$0.00 0.0029 0.050 39
GPT-Engineer 42.08 $\pm$0.00 58.78 $\pm$0.00 48.76 $\pm$0.00 0.0037 0.080 45
Self-Collab. 42.61 $\pm$0.06 55.63 $\pm$0.10 47.82 $\pm$0.08 0.0030 0.066 69
MapCoder 40.59 $\pm$0.26 57.49 $\pm$0.07 47.35 $\pm$0.18 0.0293 0.624 294
ChatDev 43.15 $\pm$0.51 57.62 $\pm$0.08 48.94 $\pm$0.34 0.0387 1.006 341
MetaGPT 0.00 $\pm$0.00 1.73 $\pm$0.06 0.69 $\pm$0.02 0.0290 0.786 151
Qwen-7B Vanilla LLM 22.37 $\pm$0.00 34.75 $\pm$0.00 27.32 $\pm$0.00 0.0003 0.005 32
GPT-Engineer 24.03 $\pm$0.00 39.94 $\pm$0.00 30.39 $\pm$0.00 0.0003 0.007 31
Self-Collab. 20.65 $\pm$1.00 33.37 $\pm$0.63 25.74 $\pm$0.85 0.0003 0.006 52
MapCoder 11.90 $\pm$1.61 26.96 $\pm$1.12 17.92 $\pm$1.41 0.0024 0.049 236
ChatDev 10.96 $\pm$1.73 21.35 $\pm$2.10 15.12 $\pm$1.88 0.0075 0.187 344
MetaGPT 0.00 $\pm$0.00 1.98 $\pm$0.28 0.79 $\pm$0.11 0.0207 0.095 301
Table 3: Benchmark analysis of effectiveness and efficiency. The grey row denotes the Vanilla baseline. Red and green indicate performance better and worse than Vanilla, respectively. Standard deviations are shown in small grey text. For effectiveness, the best result is shown in bold and the second-best is underlined.
This section enables automated benchmark analysis. E2EDev is an automated testing benchmark suite, accessible even to researchers less familiar with software testing procedures [^3] Cf. Appendix D for background knowledge of project/software testing..
@@ 4.1 Experiment Setup
Baselines and LLM Backbones. We consider a representative set of E2ESD methods, covering three categories: (1) Vanilla LLM, which directly generates an executable project via prompting; (2) Single-agent frameworks, represented by GPT-Engineer (Engineer, 2024), where a single LLM constructs the entire project in a single reasoning process; and (3) Multi-agent frameworks, including Self-Collaboration (Dong et al., 2024), MetaGPT (Hong et al., 2024), MapCoder (Islam et al., 2024), and ChatDev (Qian et al., 2024). We additionally include two widely used coding agents, OpenHands (Wang et al., 2024) and SWE-Agent (Yang et al., 2024), as reference systems, with results reported in Appendix C.1 since they target a different task setting from E2EDev. To disentangle the effects of framework design from model capacity, we run all baselines across diverse LLM backbones with varying scales and architectures (Hurst et al., 2024; Hui et al., 2024), including Claude-Haiku 4.5, GPT-4o, GPT-4o-mini, Qwen2.5-7B, Qwen2.5-70B, and the mixture-of-experts Qwen2.5-Max. Implementation details are provided in Appendix B.1.
Evaluation Metrics. For comprehensive evaluation, we assess code effectiveness (how well the generated code meets user requirements) and generation efficiency (the cost and time of code generation). Effectiveness is averaged over three runs per task, while efficiency is measured from a single run due to negligible variance. For code effectiveness, we evaluate how well a generated project satisfies requirements at two levels. (1) Req. Acc measures the fraction of requirements that are fully satisfied across a project, i.e., all test cases associated with a requirement pass; (2) Test Acc measures the fraction of all test cases that pass across a project; (3) Balanced Score, a weighted combination of Req. Acc and Test Acc to mitigate bias from varying test case granularity per requirement. Additionally, generation efficiency is assessed using the following three metrics: (1) Cost (USD), the average API pricing for LLM token usage per project; (2) Carbon Footprint (CO2), the average carbon footprint for generating one project; (3) Duration, the average wall-clock time for project generation. Metric details are presented in Appendix B.2.
@@ 4.2 Main Results
Limited effectiveness: While existing methods meet broad project requirements, their performance remains below acceptable standards due to limited proficiency in handling specifics. Off-the-shelf models excel at function-level tasks (e.g., HumanEval (Chen et al., 2021) in Appendix C.4) but underperform on E2ESD, achieving only 30%–50% Req. Acc. on the proposed E2EDev dataset (Table 3), with even advanced models like Claude-Haiku 4.5 and GPT-4o failing to exceed 60%. Furthermore, Soft Req. Acc., which allow partial fulfillment, suggest failures stem from their lack of proficiency in handling detailed functional specifics during implementation. This phenomenon, visualized in Figure 3, persists across different LLM backbones and E2ESD frameworks. The observed discrepancy of over 25% in both metrics suggests that while models can implement the required functionality, they often fail to address complex edge cases. Also, this issue is reflected in $\sim$10% performance gap observed on HumanEval and HumanEval-ET (cf. Appendix C.4).
@@
Figure 3: Soft-Requirement and Requirement Accuracy for Vanilla LLM (V-LLM), GPT-Engineer (GPT-E), and Self-Collaboration (Self-C). Additional results for other backbones are in Appendix C.2, Table 10.
[…]
Figure 4: Comparisons across agentic frameworks and the Vanilla LLM (on Req. Acc.).
Agentic frameworks do not consistently improve the effectiveness of base LLMs. While agentic frameworks offer the potential for performance gains, their inherent complexity imposes significant demands on the foundational capabilities of the underlying LLM. Figure 4 shows framework performance varies widely across LLMs, with Vanilla LLMs occasionally surpassing frameworks (and vice versa), revealing architectural overhead. Notably, multi-agent frameworks often underperform single-agent approaches due to the necessary coordination demands among agents, such as task decomposition (Xia et al., 2024; Han et al., 2024), role/instruction following (Pan et al., 2025; Hammond et al., 2025), dialogue state/context management (Pan et al., 2025; Li et al., 2023; Hammond et al., 2025), and the increasing interaction costs as shown in Table 4. Therefore, error accumulation from these complexities can degrade performance, particularly in E2ESD tasks. Overall, our results indicate that current frameworks rely heavily on strong base LLMs, suggesting that future designs should reduce this reliance to improve robustness across diverse LLM backbones.
MetaGPT suffers from communication breakdowns in its multi-agent architecture. Even at the function level, we observe evidence of communication failures on HumanEval using the official implementation (Appendix C.4), which already limits its performance. At the project level, the same issue becomes more severe, causing MetaGPT to fail on nearly all test cases and requirements, even with strong LLMs (see Section 4.3).
4.3 Failure Mode Analysis
Figure 5: Requirement-level error distribution across frameworks. Pie charts show error proportions (centers indicate total errors), and trend lines report absolute counts across frameworks.
Setup. To understand underlying failure causes, we conducted a rigorous human evaluation on 360 projects (randomly sampled from 10 entries across 6 LLMs and 6 frameworks). Four domain experts evaluate each project along four dimensions: Code Inconsistency (e.g., missing, conflicting, or empty functions), Requirement Missing (required features not implemented), Requirement Misaligned (implementation logic deviates from the requirements), and Detail Mismatch (mostly correct but with minor errors). Figure 5 presents the distribution of failure modes for existing frameworks on selected LLM backbones (GPT-4o, Qwen2.5-70B, and Qwen-Max). Detailed results and case studies are in Appendices C.2 and C.5. Human evaluation protocol and agreement are in Appendix B.3.
Model Method $_{\textbf{prompt}} Tok$$_{\textbf{compl.}}$ Turns
@@ 4.3 Failure Mode Analysis
Table 4: Statistics of interaction overhead across representative frameworks. Turns = Dialogue Turns
Analysis of Code InconsistencyVanilla LLMs generally maintain high code consistency, whereas this consistency often degrades under multi-agent settings. This degradation is largely caused by exposure to excessive or irrelevant context. For instance, MapCoder employs analogical prompting (Yasunaga et al., 2024), retrieving prior projects from memory for few-shot guidance. However, these exemplars often share only a coarse domain similarity (e.g., web applications) rather than functional relevance, introducing noise instead of actionable guidance. In contrast, frameworks that restrict coder inputs to task-specific requirements and component analyses effectively reduce interference, leading to more stable code generation.
Analysis of Requirement MissingFlawed multi-agent designs can lead to overlooked requirements. As shown in Figure 5, Vanilla LLM and ChatDev—the only methods without direct access to requirement analysis during coding [^4] GPT-Engineer implicitly performs requirement analysis via a “think step by step” process.—exhibit the most severe requirement omissions. In ChatDev, although a requirement analysis agent exists, the coding agent only receives high-level guidance from the Executive Officer, without access to detailed requirement specifications. This limited visibility substantially increases the risk of missing essential requirements, highlighting the necessity of exposing coding agents to explicit requirement analyses.
@@
Analysis of Requirement MisalignmentCore component analysis helps align code with user requirements. As shown in Figure 5, multi-agent frameworks significantly reduce requirement misalignment. This benefit mainly arises from identifying key structural and functional components, such as HTML layout and JavaScript logic, guiding the coding agent toward user-specified logic and structure. Thus, the generated code more faithfully reflects the intended functionality.
Analysis of Detail MismatchLLMs struggle with fine-grained detail control, which is exacerbated in multi-agent systems due to longer contextual chains. Figure 5 shows that detail mismatches occur across all frameworks, with Vanilla LLM exhibiting the fewest and Self-Collaboration the most. We group these errors into three categories: (1) Unhandled logical edge cases, which often span multiple modules and complex state dependencies and remain difficult for all frameworks; (2) Missing validation or error handling; and (3) UI display inconsistencies. The latter two mainly result from progressive dilution of detailed constraints as requirements pass through multiple agents that prioritize high-level functionality. Frameworks such as MapCoder and ChatDev partially mitigate this issue by repeatedly prepending original requirements and prior messages, helping preserve critical details.
Figure 6: Error distribution of MetaGPT
@@
Analysis of MetaGPTInter-agent communication breakdowns impair code consistency and requirement adherence. As shown in Figure 6, 44% of MetaGPT failures arise from code inconsistency, including missing files and syntax errors. Over 43% of these occur when the programmer agent ignores the architect’s prescribed file structure. In addition, 54% of syntax errors stem from malformed imports or unmatched tokens, likely because the engineer agent is simultaneously responsible for code generation and tool invocation, with tool-related prompts outweighing coding instructions by over 10×. Moreover, 25% of failures are caused by missing requirements, and 31% by incomplete or inconsistent implementations. This is partly due to the product manager repeatedly rewriting or compressing the original requirements, rather than restating them as instructed. Similar behavior is observed on HumanEval, where MetaGPT with GPT-4o achieves only $\sim$50% pass rate, and nearly 30% of outputs fail to match required function names. Overall, these results indicate that excessive agent decomposition and communication overhead can undermine both specification adherence and coding reliability.
5 Conclusion
Our E2EDev represents a pivotal step toward standardizing the evaluation of E2ESD frameworks by providing detailed requirement specifications and reliable evaluation protocols. While LLMs have shown strong capabilities in isolated engineering tasks, our work highlights the significant challenges remaining in achieving fully automated E2ESD, constrained by both inherent model limitations and communication inefficiencies. Consequently, future work should prioritize designing effective, reliable, and cost-efficient E2ESD solutions that can fully realize the potential of LLMs in automating software development.
Limitations